# Unbiased Regression-Adjusted Estimation of Average Treatment Effects in Randomized Controlled Trials

### v1

@@ B.2 Omitted Proofs of Section˜4
We first show that the estimator is unbiased. Let
${\boldsymbol{\Lambda}}_{\lambda}^{(-i)}=({\bm{X}}_{-i}^{\top}{\bm{X}}_{-i}+\lambda{\bm{I}})^{-1}.$
By law of total expectation, we have
@@
$\widetilde{{\bm{t}}}^{(-i)}=\begin{cases}\widetilde{{\bm{t}}}^{(1)}&\text{if }z_{i}=1,\\ \widetilde{{\bm{t}}}^{(0)}&\text{if }z_{i}=-1.\end{cases}$ (24)
@@
$\widetilde{{\bm{y}}}^{(-i)}=\frac{\widetilde{{\boldsymbol{\mu}}}}{n}+\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{{\bm{t}}}^{(-i)}}{n}.$ (26)
@@
$v_{i}z_{i}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})-\frac{1}{n}(y^{(1)}_{i}-y^{(0)}_{i})=v_{i}z_{i}(\widetilde{{\boldsymbol{\mu}}}_{i}/n-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})\,.$
@@
$\mathrm{Var}(\widehat{\tau})\;=\;{\mathbb{E}}\bigl[(\widehat{\tau}-\tau)^{2}\bigr]\;=\;{\mathbb{E}}\Biggl[\Bigl(\sum_{i=1}^{n}v_{i}\,z_{i}\,\Bigl(\tfrac{\widetilde{\boldsymbol{\mu}}_{i}}{n}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}\Bigr)\Bigr)^{2}\Biggr].$ (27)
@@
$\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\;=\;{\boldsymbol{\Lambda}}_{\lambda}^{(-i)}\,{\bm{X}}_{-i}^{\top}\,\left(\tfrac{\widetilde{\boldsymbol{\mu}}}{n}+\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{\bm{t}}^{(-i)}}{n}\right)_{-i}.$
@@
$={\mathbb{E}}\Biggl[\Biggl(\sum_{i=1}^{n}v_{i}\,z_{i}\,\Bigl\{\frac{\widetilde{\boldsymbol{\mu}}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{n\,(1-h_{\lambda ii})}-\frac{{\bm{x}}_{i}^{\top}\,\boldsymbol{\Lambda}_{\lambda}\,{\bm{X}}_{-i}^{\top}}{(1-h_{\lambda ii})}\,\left(\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{\bm{t}}^{(-i)}}{n}\right)_{-i}\Bigr\}\Biggr)^{2}\Biggr],$ (29)
where the second equality follows from Lemmas˜A.2 and B.1. For a more compact notation, we denote $\overline{h}_{i}:=(1-h_{\lambda ii})$ for the rest of this proof. We decompose (29) into three components and compute the corresponding expectations separately.
@@
    $T_{1}={\mathbb{E}}\left[\sum_{i,j\in[n]}\frac{v_{i}\,z_{i}\,v_{j}\,z_{j}\,}{n^{2}}\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{i}}\,\frac{\widetilde{\mu}_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{j}}\right]$.
@@
    $T_{2}=-2\,{\mathbb{E}}\left[\sum_{i,j\in[n]}\frac{v_{i}\,z_{i}\,v_{j}\,z_{j}\,}{n^{2}}\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{i}}\;\left({\bm{x}}_{j}^{\top}\,\boldsymbol{\Lambda}_{\lambda}\,{\bm{X}}_{-j}^{\top}\,({\bm{v}}^{(-j)}{\bm{z}}\widetilde{\bm{t}}^{(-j)})_{-j}\right)\right]$.
@@
$T_{1}={\mathbb{E}}\Biggl[\Biggl(\sum_{i=1}^{n}\frac{v_{i}\,z_{i}\,}{n}\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{i}}\Biggr)^{2}\Biggr].$
We denote $R_{i}:=\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{i}}$. Since $\widetilde{{\boldsymbol{\mu}}}$ and $\boldsymbol{\beta}_{\lambda}$ are fixed vectors and do not depend on the treatment assignment vector $\bm{z}$}, by linearity of expectation,
$T_{1}=\frac{1}{n^{2}}{\mathbb{E}}\Biggl[\Biggl(\sum_{i=1}^{n}v_{i}z_{i}R_{i}\Biggr)^{2}\Biggr]=\frac{1}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}R_{i}R_{j}{\mathbb{E}}[v_{i}z_{i}v_{j}z_{j}].$
For $i=j$, ${\mathbb{E}}[(v_{i}z_{i})^{2}]={\mathbb{E}}[v_{i}^{2}]$. Therefore by defnition,
@@
$T_{1}$ $=\frac{1}{n^{2}}\left(\sum_{i=1}^{n}R_{i}^{2}{\mathbb{E}}[v_{i}^{2}]+\sum_{i\neq j}R_{i}R_{j}{\mathbb{E}}[v_{i}z_{i}v_{j}z_{j}]\right)$
$=\frac{1}{n^{2}n_{T}n_{C}}\left(\sum_{i=1}^{n}R_{i}^{2}-\frac{1}{n-1}\sum_{i,j\in[n]:i\neq j}R_{i}R_{j}\right).$
@@
$T_{1}$ $=\frac{1}{n(n-1)n_{T}n_{C}}\sum_{i=1}^{n}\left(R_{i}-\frac{1}{n}\sum_{j=1}^{n}R_{j}\right)^{2}$
$=\frac{1}{n(n-1)n_{T}n_{C}}\sum_{i=1}^{n}\left(\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{(1-h_{\lambda ii})}-\frac{1}{n}\sum_{j=1}^{n}\frac{\widetilde{\mu}_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda}}{(1-h_{\lambda jj})}\right)^{2}.$
@@
$T_{2}$ $=-2\,{\mathbb{E}}\left[\Biggl(\sum_{i=1}^{n}\frac{v_{i}z_{i}R_{i}}{n}\Biggr)\Biggl(\sum_{j=1}^{n}\frac{v_{j}z_{j}}{n\overline{h}_{j}}{\bm{x}}_{j}^{\top}{\boldsymbol{\Lambda}}_{\lambda}{\bm{X}}_{-j}^{\top}({\bm{v}}^{(-j)}{\bm{z}}\widetilde{{\bm{t}}}^{(-j)})_{-j}\Biggr)\right]$
$=\frac{-2}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{R_{i}}{\overline{h}_{j}}\sum_{k\in[n]:k\neq j}h_{\lambda jk}\,{\mathbb{E}}\left[v_{i}z_{i}v_{j}z_{j}v_{k}^{(-j)}z_{k}\widetilde{t}_{k}^{(-j)}\right]$
We denote ${\mathbb{E}}\left[v_{i}z_{i}v_{j}z_{j}v_{k}^{(-j)}z_{k}\widetilde{t}_{k}^{(-j)}\right]$. One can easily check that $E_{iik}=0$ for $i\neq k$. Therefore we consider two cases: 1) $i=k$, and 2) $i\neq k$, $i\neq j$—note that the above expression guarantees $j\neq k$. We have
$E_{iji}={\mathbb{E}}\left[v_{i}v^{(-j)}_{i}v_{j}z_{j}\widetilde{t}_{i}^{(-j)}\right]=\frac{\widetilde{\mu}_{i}}{(n-1)n_{T}n_{C}}.$ (30)
@@
$E_{ijk}={\mathbb{E}}\left[v_{i}z_{i}v_{j}z_{j}v_{k}^{(-j)}z_{k}\widetilde{t}_{k}^{(-j)}\right]=\frac{-\widetilde{\mu}_{k}}{(n-1)(n-2)n_{T}n_{C}}.$
@@
$T_{2,A}:=\frac{-2}{n^{2}(n-1)n_{T}n_{C}}\sum_{i,j\in[n]:i\neq j}\frac{R_{i}h_{\lambda ij}\widetilde{\mu}_{i}}{\overline{h}_{j}}.$
@@
$T_{2,B}=\frac{2}{n^{2}(n-1)(n-2)n_{T}n_{C}}\sum_{i,j\in[n]:i\neq j}\sum_{k\in[n]:k\neq i,j}\frac{R_{i}h_{\lambda jk}\widetilde{\mu}_{k}}{\overline{h}_{j}}.$
@@
$T_{2}=T_{2,A}+T_{2,B}=\frac{-2}{n^{2}(n-1)n_{T}n_{C}}\left(\sum_{i,j\in[n]:i\neq j}\frac{R_{i}h_{\lambda ij}\widetilde{\mu}_{i}}{\overline{h}_{j}}-\frac{1}{n-2}\sum_{k\in[n]:k\neq i,j}\frac{R_{i}h_{\lambda jk}\widetilde{\mu}_{k}}{\overline{h}_{j}}\right).$
@@
$T_{3}$ $=\frac{1}{F}\sum_{i\in[n]}\biggl[\sum_{k\in[n]:k\neq i}h_{\lambda ik}^{2}\frac{(n_{C}-1)(\widetilde{\bm{t}}^{(1)}_{k})^{2}+(n_{T}-1)(\widetilde{\bm{t}}^{(0)}_{k})^{2}}{\overline{h}_{i}^{2}}$ (31)
$-\frac{1}{n-2}\sum_{l\in[n]:l\neq i,k}\frac{h_{\lambda ik}h_{\lambda il}}{\overline{h}_{i}^{2}}((n_{C}-1)\widetilde{\bm{t}}^{(1)}_{k}\widetilde{\bm{t}}^{(1)}_{l}+(n_{T}-1)\widetilde{\bm{t}}^{(0)}_{k}\widetilde{\bm{t}}^{(0)}_{l})\biggr$] (32)
$+\frac{1}{F}\sum_{i,j\in[n]:i\neq j}\sum_{k\in[n]:k\neq i,j}\frac{h_{\lambda ik}h_{\lambda jk}}{\overline{h}_{i}\overline{h}_{j}}\Bigl[a_{T}(\widetilde{\bm{t}}^{(1)}_{k})^{2}-2n\widetilde{\bm{t}}^{(1)}_{k}\widetilde{\bm{t}}^{(0)}_{k}+a_{C}(\widetilde{\bm{t}}^{(0)}_{k})^{2}\Bigr]$ (33)
$+\frac{1}{n^{3}(n-1)}\sum_{i,j\in[n]:i\neq j}\biggl[\frac{h_{\lambda ij}^{2}}{\overline{h}_{i}\overline{h}_{j}}\Bigl[\frac{\widetilde{\bm{t}}^{(1)}_{i}\widetilde{\bm{t}}^{(1)}_{j}}{n_{T}(n_{T}-1)}+\frac{\widetilde{\bm{t}}^{(0)}_{i}\widetilde{\bm{t}}^{(0)}_{j}}{n_{C}(n_{C}-1)}+\frac{\widetilde{\bm{t}}^{(1)}_{i}\widetilde{\bm{t}}^{(0)}_{j}+\widetilde{\bm{t}}^{(0)}_{i}\widetilde{\bm{t}}^{(1)}_{j}}{n_{C}n_{T}}\Bigr]$ (34)
$-\frac{1}{n-2}\sum_{k\in[n]:k\neq i,j}\frac{h_{\lambda ji}}{\overline{h}_{i}\overline{h}_{j}}\Bigl[\frac{(h_{\lambda ik}\widetilde{\bm{t}}^{(1)}_{i}+h_{\lambda jk}\widetilde{\bm{t}}^{(1)}_{j})\widetilde{\bm{t}}^{(1)}_{k}}{n_{T}(n_{T}-1)}+\frac{(h_{\lambda ik}\widetilde{\bm{t}}^{(0)}_{i}+h_{\lambda jk}\widetilde{\bm{t}}^{(0)}_{j})\widetilde{\bm{t}}^{(0)}_{k}}{n_{C}(n_{C}-1)}$ (35)
$+\frac{(h_{\lambda ik}\widetilde{\bm{t}}^{(1)}_{i}+h_{\lambda jk}\widetilde{\bm{t}}^{(1)}_{j})\widetilde{\bm{t}}^{(0)}_{k}+(h_{\lambda ik}\widetilde{\bm{t}}^{(0)}_{i}+h_{\lambda jk}\widetilde{\bm{t}}^{(0)}_{j})\widetilde{\bm{t}}^{(1)}_{k}}{n_{C}n_{T}}\Bigr]$ (36)
@@
Proof of Lemma˜6.
We argue in three steps.
Step 1: Existence and boundedness of $\boldsymbol{\beta}^{*}_{\lambda}$.
The first-order condition for the minimizer in (16) implies
$\boldsymbol{\beta}^{*}_{\lambda}\;=\;\lim_{n\to\infty}({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\frac{\widetilde{{\boldsymbol{\mu}}}}{n}.$
By Assumption˜2, $n^{-1}{\bm{X}}^{\top}{\bm{X}}\to\Sigma_{X}$, where $\Sigma_{X}$ is bounded and invertible. Therefore
$\big\|({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}\big\|_{2}\;=\;O(n^{-1}).$ (38)
Moreover,
${\bm{X}}^{\top}\frac{\widetilde{{\boldsymbol{\mu}}}}{n}=\frac{1}{n}\sum_{j=1}^{n}{\bm{x}}_{j}\,\Big(n_{C}y_{j}^{(1)}+n_{T}y_{j}^{(0)}\Big).$
By Assumptions˜2 and 3, the averages $n^{-1}\sum_{j}{\bm{x}}_{j}y_{j}^{(1)}$ and $n^{-1}\sum_{j}{\bm{x}}_{j}y_{j}^{(0)}$ converge to finite limits, and $n_{T}/n$ and $n_{C}/n$ are bounded away from 0 and $1$. Hence $\big\|{\bm{X}}^{\top}\frac{\widetilde{{\boldsymbol{\mu}}}}{n}\big\|_{2}=O(n)$. Combining this with (38) shows that
$({\bm{X}}^{\top}{\bm{X}}+\lambda{\bm{I}})^{-1}{\bm{X}}^{\top}\frac{\widetilde{{\boldsymbol{\mu}}}}{n}$
is $O(1)$ and converges to a finite limit. Thus $\boldsymbol{\beta}^{*}_{\lambda}$ exists and is bounded.
Step 2: Convergence of an auxiliary full-sample ridge coefficient.
For each $i\in[n]$, define the auxiliary (full-sample) ridge coefficient

### v2

@@ B.2 Omitted Proofs of Section˜4
We first show that the estimator is unbiased. Let ${\boldsymbol{\Lambda}}_{\lambda}^{(-i)}=({\bm{X}}_{-i}^{\top}{\bm{X}}_{-i}+\lambda{\bm{I}})^{-1}$. By law of total expectation, we have
@@
$\widetilde{{\bm{t}}}^{(-i)}=\begin{cases}\widetilde{{\bm{t}}}^{(1)}&\text{if }z_{i}=1,\\ \widetilde{{\bm{t}}}^{(0)}&\text{if }z_{i}=-1.\end{cases}$
@@
$\widetilde{{\bm{y}}}^{(-i)}=\frac{\sqrt{n_{C}n_{T}}}{n}\widetilde{{\boldsymbol{\mu}}}+\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{{\bm{t}}}^{(-i)}}{n}.$ (23)
@@
$v_{i}z_{i}(y_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})-\frac{1}{n}(y^{(1)}_{i}-y^{(0)}_{i})=v_{i}z_{i}(\frac{\sqrt{n_{C}n_{T}}}{n}\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)})\,.$
@@
${\rm Var}(\widehat{\tau})\;=\;{\mathbb{E}}\bigl[(\widehat{\tau}-\tau)^{2}\bigr]\;=\;{\mathbb{E}}\Biggl[\Bigl(\sum_{i=1}^{n}v_{i}\,z_{i}\,\Bigl(\frac{\sqrt{n_{C}n_{T}}}{n}\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\widehat{\boldsymbol{\beta}}^{(-i)}\Bigr)\Bigr)^{2}\Biggr].$ (24)
@@
$\widehat{\boldsymbol{\beta}}_{\lambda}^{(-i)}\;=\;{\boldsymbol{\Lambda}}_{\lambda}^{(-i)}\,{\bm{X}}_{-i}^{\top}\,\left(\frac{\sqrt{n_{C}n_{T}}}{n}\widetilde{{\boldsymbol{\mu}}}+\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{\bm{t}}^{(-i)}}{n}\right)_{-i}.$
@@
$={\mathbb{E}}\Biggl[\Biggl(\sum_{i=1}^{n}v_{i}\,z_{i}\,\Bigl\{\frac{\sqrt{n_{C}n_{T}}}{n}\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{(1-h_{\lambda ii})}-\frac{{\bm{x}}_{i}^{\top}\,\boldsymbol{\Lambda}_{\lambda}\,{\bm{X}}_{-i}^{\top}}{(1-h_{\lambda ii})}\,\left(\frac{{\bm{v}}^{(-i)}{\bm{z}}\widetilde{\bm{t}}^{(-i)}}{n}\right)_{-i}\Bigr\}\Biggr)^{2}\Biggr],$ (25)
where the second equality follows from Lemmas˜3 and 6. For a more compact notation, we denote $\overline{h}_{i}=(1-h_{\lambda ii})$ for the rest of this proof. We decompose (25) into three components and compute the corresponding expectations separately.
@@
    $T_{1}={\mathbb{E}}\left[\sum_{i,j\in[n]}\frac{(n_{C}n_{T})v_{i}\,z_{i}\,v_{j}\,z_{j}\,}{n^{2}}\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{i}}\,\frac{\widetilde{\mu}_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{j}}\right]$.
@@
    $T_{2}=-2\,{\mathbb{E}}\left[\sum_{i,j\in[n]}\frac{\sqrt{n_{C}n_{T}}v_{i}\,z_{i}\,v_{j}\,z_{j}\,}{n^{2}}\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{i}\overline{h}_{j}}\;\left({\bm{x}}_{j}^{\top}\,\boldsymbol{\Lambda}_{\lambda}\,{\bm{X}}_{-j}^{\top}\,({\bm{v}}^{(-j)}{\bm{z}}\widetilde{\bm{t}}^{(-j)})_{-j}\right)\right]$.
@@
$T_{1}={\mathbb{E}}\Biggl[\Biggl(\sum_{i=1}^{n}\frac{\sqrt{n_{C}n_{T}}v_{i}\,z_{i}\,}{n}\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{i}}\Biggr)^{2}\Biggr].$
We denote $R_{i}=\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{\overline{h}_{i}}$. Since $\widetilde{{\boldsymbol{\mu}}}$ and $\boldsymbol{\beta}_{\lambda}$ are fixed vectors and do not depend on the treatment assignment vector $\bm{z}$}, by linearity of expectation,
$T_{1}=\frac{n_{C}n_{T}}{n^{2}}{\mathbb{E}}\Biggl[\Biggl(\sum_{i=1}^{n}v_{i}z_{i}R_{i}\Biggr)^{2}\Biggr]=\frac{n_{C}n_{T}}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}R_{i}R_{j}{\mathbb{E}}[v_{i}z_{i}v_{j}z_{j}].$
For $i=j$, ${\mathbb{E}}[(v_{i}z_{i})^{2}]={\mathbb{E}}[v_{i}^{2}]$. Therefore by definition,
@@
$T_{1}$ $=\frac{n_{T}n_{C}}{n^{2}}\left(\sum_{i=1}^{n}R_{i}^{2}{\mathbb{E}}[v_{i}^{2}]+\sum_{i\neq j}R_{i}R_{j}{\mathbb{E}}[v_{i}z_{i}v_{j}z_{j}]\right)$
$=\frac{1}{n^{2}}\left(\sum_{i=1}^{n}R_{i}^{2}-\frac{1}{n-1}\sum_{i,j\in[n]:i\neq j}R_{i}R_{j}\right).$
@@
$T_{1}$ $=\frac{1}{n(n-1)}\sum_{i=1}^{n}\left(R_{i}-\frac{1}{n}\sum_{j=1}^{n}R_{j}\right)^{2}$
$=\frac{1}{n(n-1)}\sum_{i=1}^{n}\left(\frac{\widetilde{\mu}_{i}-{\bm{x}}_{i}^{\top}\boldsymbol{\beta}_{\lambda}}{(1-h_{\lambda ii})}-\frac{1}{n}\sum_{j=1}^{n}\frac{\widetilde{\mu}_{j}-{\bm{x}}_{j}^{\top}\boldsymbol{\beta}_{\lambda}}{(1-h_{\lambda jj})}\right)^{2}.$
@@
$T_{2}$ $=-2\,{\mathbb{E}}\left[\Biggl(\sum_{i=1}^{n}\frac{\sqrt{n_{T}n_{C}}v_{i}z_{i}R_{i}}{n}\Biggr)\Biggl(\sum_{j=1}^{n}\frac{v_{j}z_{j}}{n\overline{h}_{j}}{\bm{x}}_{j}^{\top}{\boldsymbol{\Lambda}}_{\lambda}{\bm{X}}_{-j}^{\top}({\bm{v}}^{(-j)}{\bm{z}}\widetilde{{\bm{t}}}^{(-j)})_{-j}\Biggr)\right]$
$=\frac{-2\sqrt{n_{T}n_{C}}}{n^{2}}\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{R_{i}}{\overline{h}_{j}}\sum_{k\in[n]:k\neq j}h_{\lambda jk}\,{\mathbb{E}}\left[v_{i}z_{i}v_{j}z_{j}v_{k}^{(-j)}z_{k}\widetilde{t}_{k}^{(-j)}\right]$
We denote $E_{ijk}:={\mathbb{E}}\left[v_{i}z_{i}v_{j}z_{j}v_{k}^{(-j)}z_{k}\widetilde{t}_{k}^{(-j)}\right]$. One can easily check that $E_{iik}=0$ for $i\neq k$. Therefore we consider two cases: 1) $i=k$, and 2) $i\neq k$, $i\neq j$—note that the above expression guarantees $j\neq k$. We have
$E_{iji}={\mathbb{E}}\left[v_{i}v^{(-j)}_{i}v_{j}z_{j}\widetilde{t}_{i}^{(-j)}\right]=\frac{\widetilde{\mu}_{i}}{(n-1)\sqrt{n_{T}n_{C}}}.$ (26)
@@
$E_{ijk}={\mathbb{E}}\left[v_{i}z_{i}v_{j}z_{j}v_{k}^{(-j)}z_{k}\widetilde{t}_{k}^{(-j)}\right]=\frac{-\widetilde{\mu}_{k}}{(n-1)(n-2)\sqrt{n_{T}n_{C}}}.$
@@
$T_{2,A}=\frac{-2}{n^{2}(n-1)}\sum_{i,j\in[n]:i\neq j}\frac{R_{i}h_{\lambda ij}\widetilde{\mu}_{i}}{\overline{h}_{j}}.$
@@
$T_{2,B}=\frac{2}{n^{2}(n-1)(n-2)}\sum_{i,j\in[n]:i\neq j}\sum_{k\in[n]:k\neq i,j}\frac{R_{i}h_{\lambda jk}\widetilde{\mu}_{k}}{\overline{h}_{j}}.$
@@
$T_{2}=T_{2,A}+T_{2,B}=\frac{-2}{n^{2}(n-1)}\sum_{i,j\in[n]:i\neq j}\left(\frac{R_{i}h_{\lambda ij}\widetilde{\mu}_{i}}{\overline{h}_{j}}-\frac{1}{n-2}\sum_{k\in[n]:k\neq i,j}\frac{R_{i}h_{\lambda jk}\widetilde{\mu}_{k}}{\overline{h}_{j}}\right).$
@@
$T_{3}$ $=\frac{1}{F}\sum_{i\in[n]}\sum_{k\in[n]:k\neq i}\biggl[h_{\lambda ik}^{2}\frac{(n_{C}-1)(\widetilde{\bm{t}}^{(1)}_{k})^{2}+(n_{T}-1)(\widetilde{\bm{t}}^{(0)}_{k})^{2}}{\overline{h}_{i}^{2}}$
$-\frac{1}{n-2}\sum_{l\in[n]:l\neq i,k}\frac{h_{\lambda ik}h_{\lambda il}}{\overline{h}_{i}^{2}}((n_{C}-1)\widetilde{\bm{t}}^{(1)}_{k}\widetilde{\bm{t}}^{(1)}_{l}+(n_{T}-1)\widetilde{\bm{t}}^{(0)}_{k}\widetilde{\bm{t}}^{(0)}_{l})\biggr$]
$+\frac{1}{F}\sum_{i,j\in[n]:i\neq j}\sum_{k\in[n]:k\neq i,j}\frac{h_{\lambda ik}h_{\lambda jk}}{\overline{h}_{i}\overline{h}_{j}}\Bigl[a_{T}(\widetilde{\bm{t}}^{(1)}_{k})^{2}-2n\widetilde{\bm{t}}^{(1)}_{k}\widetilde{\bm{t}}^{(0)}_{k}+a_{C}(\widetilde{\bm{t}}^{(0)}_{k})^{2}\Bigr]$
$+\frac{1}{n^{3}(n-1)}\sum_{i,j\in[n]:i\neq j}\biggl[\frac{h_{\lambda ij}^{2}}{\overline{h}_{i}\overline{h}_{j}}\Bigl[\frac{\widetilde{\bm{t}}^{(1)}_{i}\widetilde{\bm{t}}^{(1)}_{j}}{n_{T}(n_{T}-1)}+\frac{\widetilde{\bm{t}}^{(0)}_{i}\widetilde{\bm{t}}^{(0)}_{j}}{n_{C}(n_{C}-1)}+\frac{\widetilde{\bm{t}}^{(1)}_{i}\widetilde{\bm{t}}^{(0)}_{j}+\widetilde{\bm{t}}^{(0)}_{i}\widetilde{\bm{t}}^{(1)}_{j}}{n_{C}n_{T}}\Bigr]$
$-\frac{1}{n-2}\sum_{k\in[n]:k\neq i,j}\frac{h_{\lambda ji}}{\overline{h}_{i}\overline{h}_{j}}\Bigl[\frac{(h_{\lambda ik}\widetilde{\bm{t}}^{(1)}_{i}+h_{\lambda jk}\widetilde{\bm{t}}^{(1)}_{j})\widetilde{\bm{t}}^{(1)}_{k}}{n_{T}(n_{T}-1)}+\frac{(h_{\lambda ik}\widetilde{\bm{t}}^{(0)}_{i}+h_{\lambda jk}\widetilde{\bm{t}}^{(0)}_{j})\widetilde{\bm{t}}^{(0)}_{k}}{n_{C}(n_{C}-1)}$
$+\frac{(h_{\lambda ik}\widetilde{\bm{t}}^{(1)}_{i}+h_{\lambda jk}\widetilde{\bm{t}}^{(1)}_{j})\widetilde{\bm{t}}^{(0)}_{k}+(h_{\lambda ik}\widetilde{\bm{t}}^{(0)}_{i}+h_{\lambda jk}\widetilde{\bm{t}}^{(0)}_{j})\widetilde{\bm{t}}^{(1)}_{k}}{n_{C}n_{T}}\Bigr]$
@@
Proof of Theorem˜6.
We define: